OpenStudy (anonymous):

A curved road of diameter 2 Km is banked so no friction acts at maximum speed of 30m/s what is the banked angle?

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

\[\tan \theta = V^2/rg\]

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

@experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX @experimentX

OpenStudy (experimentx):

seems all right ...

OpenStudy (anonymous):

\[\tan \theta = 0.09\]

OpenStudy (anonymous):

is that correct

OpenStudy (experimentx):

begin with figure ..

OpenStudy (anonymous):

@experimentX wat answer u got

OpenStudy (experimentx):

haven't got any ... :((

OpenStudy (anonymous):

ok...pplzz solve it

OpenStudy (experimentx):